The line of which equation below is perpendicular to y=5/8x–2?
A) y=5/8x+5 B) y=-5/8x–2 C) y=-8/5x+5 D) y=8/5x–4

Answer by ewatrrr(24785)


Hi
referring to the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
y= 5/8x –2 has a slope m = (5/8). A line ⊥ to it would have a slope m = -8/5
as ⊥ lines have slopes that are 'negative reciprocals' of one another